OpenStudy (anonymous):

What is the simplified expression for 3^3 • 3^3 over 3^4

OpenStudy (anonymous):

I think it would be D @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you forgot to list the choices

OpenStudy (anonymous):

3^0 3^1 3^2 3^4

jimthompson5910 (jim_thompson5910):

3^3 times 3^3 = 3^(3+3) = 3^6 then 3^6 over 3^4 = 3^(6-4) = 3^2

OpenStudy (anonymous):

So it would be 3^4

jimthompson5910 (jim_thompson5910):

Rules \[\Large a^b*a^c = a^{b+c}\] \[\Large \frac{a^b}{a^c} = a^{b-c}\]

jimthompson5910 (jim_thompson5910):

no, at the end I got 3^2

jimthompson5910 (jim_thompson5910):

hopefully you see how/why

OpenStudy (anonymous):

I do now, thank you.

jimthompson5910 (jim_thompson5910):

np